Use the traditional method to test the given hypothesis. Assume that the population is normally
distributed and that the sample has been randomly selected. A manufacturer uses a new production
method to produce steel rods. A random sample of 37 steel rods resulted in lengths with a mean 6 and
standard deviation of 4.7 cm. At the 0.10 significance level, test the claim that the new production method
has mean 5.5 cm, which was the mean for the old method.
Answer #1

S = 4.1 (10) The test statstic t²9-M (6-5.5) Significance level n=37 (i) Ho: M=55 (11) a=010 Hai u&55 islo t=[ooburi at=36 (1